OpenStudy (anonymous):

Hi everyone, I need some help with my Algebra homework. Any help is appreciated :) Find the equation in standard form of the parabola passing through the points. (3,-6), (1, -2), (6,3) For the second question it is the same thing. The points are (-1, -1), (1,3), and (-3,1) I am having a really hard time understanding how to do this. Thank you

OpenStudy (anonymous):

Thank you, can you explain your steps to me so I can learn to do this on my own?

OpenStudy (anonymous):

Is there anyone else able/available to help me through these questions?

OpenStudy (amistre64):

with 3 random points, one method would be to define 3 equations in 3 unknowns

OpenStudy (amistre64):

a parabola takes the form: ax^2 + bx + c = y we know the x and y parts, it is the abc parts that are unknown to us

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

(3,-6), (1, -2), (6,3) when x=3 and y=-6 a(3)^2 + b(3) + c = -6 9a + 3b + c = -6 --------------------------- when x=1 and y=-2 a(1)^2 + b(1) + c = -2 a + b + c = -2 --------------------------- when x=6 and y=3 a(6)^2 + b(6) + c = 3 36a + 6b + c = 3 --------------------------- this forms a system of 3 equations; in 3 unknowns 9a + 3b + c = -6 a + b + c = -2 36a +6b + c = 3

OpenStudy (amistre64):

there are different methods that can be used to find the abc values .. substitution, elimination, and matrix methods tend to be used the most

OpenStudy (anonymous):

Ok, so if we use elimination what would happen next?

OpenStudy (amistre64):

you would use the elimination method then ... are you familiar with it and just unsure how to use it for a 3 equation setup?

OpenStudy (anonymous):

I am familiar with it, just a little bit confused.

OpenStudy (amistre64):

we wold use one of the equations to eliminate a variable out of the other 2 .. instead of just out of "the other one". Since the middle equation is fairly naked, it looks like a good one to use: scale it by some value k, to eliminate an unknown ka + kb +kc = -2k 9a + 3b + c = -6 scale it by some value n, to eliminate the same unknown in the other equation na + nb + nc = -2n 36a +6b + c = 3 i spose the question now is, which variable do you want to eliminate first?

OpenStudy (anonymous):

I guess c

OpenStudy (amistre64):

then id say let k = -1, and let n = -1 and add; what equations do we end up with?

OpenStudy (anonymous):

I am a little bit confused with the way you set up the equations for elimination. I'm sorry

OpenStudy (amistre64):

-a - b - c = 2 9a +3b + c = -6 ------------------- 8a +2b = -4 -a - b - c = 2 36a +6b + c = 3 ------------------ 35a +5b = 5

OpenStudy (amistre64):

i simply picked one of the equations to modify in order to eliminate a variable from the other 2

OpenStudy (amistre64):

that is all we did with a system of 2 equations in 2 unkowns; pick one and modify it to eliminate a variable in the other equation

OpenStudy (anonymous):

Ok, it is starting to make more sense

OpenStudy (amistre64):

the elimination we just did creates 2 equations in 2 unknowns .. which can be worked again 8a +2b = -4 35a +5b = 5 this should be something you are more used to dealing with

OpenStudy (anonymous):

Yes, so that answer is 43a+7b = 1

OpenStudy (amistre64):

no, we want to run another elimination on the system we created. im going to reduce the coefficients to make the numbers smaller 8a +2b = -4 reduce by dividing off 2 35a +5b = 5 reduce by dividing off 5 4a +b = -2 7a +b = 1 b looks like an easy target to eliminate; multiply the first equation by -1 -4a -b = 2 7a +b = 1 ------------- 3a = 3 solving for a is simple now; a = 1

OpenStudy (anonymous):

Oh, Ok. Got it.

OpenStudy (amistre64):

now that we know a, take an ab setup to solve for b 7a +b = 1 7 +b = 1 b = -6 a = 1, b = -6 now that we know ab, take an abc setup to solve for c a + b + c = -2 1 - 6 + c = -2 -5 + c = -2 c = 3 we now know all the abc parts to fill in the quadratic: ax^2 + bx + c = y x^2 -6x +3 = y

OpenStudy (anonymous):

Yea! thank you so much for helping me! It makes a lot more sense now

OpenStudy (amistre64):

youre welcome; that is what we are spose to be here for instead of like that first poster did ... just giving out an answer has no learning value :/

OpenStudy (anonymous):

No, I need to be able to work things out myself. Thank you for helping :)

OpenStudy (amistre64):

i have class starting soon so i have to go, good luck with this

OpenStudy (anonymous):

Have a great day